In late August, a particular star is seen to rise above the eastern horizon at 9 pm. Tonight (approximately two months later), the same star will rise at approximately what hour?
Answer #1

A star rises 4 minutes early from previous day.i.e 2 hours early every month, so two months later the star will rise at

9 - (2x2) =9-4 = 5PM

the same star will rise at 5 pm after two months.

Please Upvote If you find my answer Helpful.

Thank you